1. a. Susie’s annual salary is $90,000. Each month the following items are deducted from her paycheck

Federal income tax

800

Health insurance premium

200

State income tax

300

Group term life insurance premium

75

Social security tax

465

Parking

120

Medicare tax

108

Union due

50

Her gross income is $______; her disposable income is $______.

b. Based on Susie’s cash flow statement, Susie’s total inflow is $95,000, and her total outflow is $100,000, she has a _______ (surplus or deficit) of $______.
Answer #1

Her Gross income is $7500 ( Annual income/12=90000/12); her disposable income is $5382 (annual disposable Income/12=64584/12)

Annual disposable income =Annual income-Sum of all monthly deduction*12=90000-2118*12=64548

Based on Susie's cash flow statement, Susie's total inflow is $95,000, and her total outflow is $100,000, she has a deficit of $5,000